Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Squash- weather inclement

Expert replies
Source: — Problem Solving |

by anshumishra » Thu Dec 23, 2010 3:43 pm
[email protected] wrote:I don't agree with Manhattan's reasoning.

According to them the correct answer is 32.5%.

Please explain.
Image
Probability that I'll play squash if (My mother doesn't come Tomorrow) AND (Doesn't hail) AND (No Rain)
= 1/2*3/4*9/10 = 27/80 ~= 33%
Join the discussion

by [email protected] » Thu Dec 23, 2010 5:01 pm
Unfortunately, that's not the correct answer.
Join the discussion

by goyalsau » Thu Dec 23, 2010 5:48 pm
[email protected] wrote:Unfortunately, that's not the correct answer.

HI! Archit, Different kind of a question,

Its very rare that anshu got this one wrong.... after all the post that i have seen of him,

According to me, Answer should be 32.5%

There 25% of Hail & 10% chance of rain.

So 25 + 10 = 35 %

100 - 35 = 65 % chance of clear whether.

65/2 = 32.5% Mother come to bother him.

remaining 32.5% chance of play...
Saurabh Goyal
[email protected]
-------------------------


EveryBody Wants to Win But Nobody wants to prepare for Win.
Join the discussion

by anshumishra » Thu Dec 23, 2010 5:56 pm
goyalsau wrote:
[email protected] wrote:Unfortunately, that's not the correct answer.

HI! Archit, Different kind of a question,

Its very rare that anshu got this one wrong.... after all the post that i have seen of him,

According to me, Answer should be 32.5%

There 25% of Hail & 10% chance of rain.

So 25 + 10 = 35 %

100 - 35 = 65 % chance of clear whether.

65/2 = 32.5% Mother come to bother him.

remaining 32.5% chance of play...
Thanks goyalsau for the kind words and the solution !
I guess the word "hail" didn't make sense to me and now I got it that it points to a weather condition which is not favorable.
Your solution makes perfect sense to me.

Thanks
Join the discussion

by captcha » Thu Dec 23, 2010 7:09 pm
mother come = 50% chance

Hail = 25% chance

Rain = 10% chance

You can play if mom doesn't come & no hail & no rail

Probability of playing = P(no mom)*P(no hail)*P(no rain)

P(play) = .5*.75*.9

P(play)=33.75%

if hail & rain cannot come together;

P(play) = .5*(1-.25-.1)

P(play) = 32.5%
Join the discussion